OpenStudy (anonymous):

Solve x^2 + 5x + 6 = 0 by completing the square.

OpenStudy (anonymous):

\[x^2+5x+6=0\\ \left(x+\tfrac{5}{2}\right)^2-\tfrac{25}{4}+6=0\\ \left(x+\tfrac{5}{2}\right)^2=\tfrac{1}{4}\\ x+\tfrac{5}{2}=\pm\tfrac{1}{2}\\ x=-2,-3 \]